OpenStudy (anonymous):

use elementary row operations to transform each augmented coefficient matrix to echelon form. Then solve the system by back substitution. 2x_1+5x_2+12x_3=6 3x_1+x_2+5x_3=12 5x_1+8x_2+21x_3=17

OpenStudy (anonymous):

2 5 12 6 3 1 5 12 5 8 21 17 that's how I put it in the matrix. Then I did -3/2(R1+R2) and got.. 2 5 12 6 0 -13/2 -13 3 5 8 21 17 then I did.. -5/2(R1+R3) and got.. 2 5 12 6 0 -13/2 -13 3 0 -9/2 -9 2

OpenStudy (anonymous):

@UnkleRhaukus can you help me? :) please

OpenStudy (unklerhaukus):

why did you use those row operations?

OpenStudy (unklerhaukus):

PS i dont think this system has a solution

OpenStudy (anonymous):

I'm not familiar with this topic.. so im just trying things out.. I'm not sure how to choose the 'right' row operations. how do you know if it doesnt have a solution?

OpenStudy (unklerhaukus):

i have a matrix solving program that is helpful sometimes. .

OpenStudy (unklerhaukus):

*** lets try some different row operations (LESS mistakes this time) 2 5 12 6 3 1 5 12 5 8 21 17 2 5 12 6 1 -4* -7 6 R_2 becomes R_2-R_1 1 -2 -3 5 R_3 becomes R_3-2R_1

OpenStudy (unklerhaukus):

now 0 9 18 -4 R_1 becomes R_1-2R_3 0 -6* -10* 1 R_2 becomes R_2-R_3 1 -2 -3 5

OpenStudy (unklerhaukus):

1 -2 -3 5 R_1 becomes R_3 0 1 5/3 -1/6 R_2 becomes R_2/-6 0 9 18 -4 R_3 becomes R_1 1 -2 -3 5 0 1 5/3 -1/6 0 1 2 -4/9 R_3 becomes R_3/9 1 -2 -3 5 0 1 5/3 -1/6 0 0 1/3 -5/18 R_3 becomes R_3-R_2 1 -2 -3 5 0 1 5/3 -1/6 0 0 1 -5/6 R_3 becomes 3*R_3

OpenStudy (unklerhaukus):

1 0 1/3 14/3 R_1 becomes R_1+2R_2 0 1 0 11/9 R_2 becomes R_2-5/3 R_3 0 0 1 -5/6 R_3 becomes 3*R_3 1 0 0 89/18 R_1 becomes R_1-1/3 R_3 0 1 0 11/9 0 0 1 -5/6 lets see if this works ___ 2x_1+5x_2+12x_3=6 2(89/18)+5(11/9)+12(-5/6)= 6 \(\checkmark\) ___ 3x_1+x_2+5x_3=12 3(89/9)+11/9+5(-5/6) ~ 11.888... ≠12 (not right) ___ 5x_1+8x_2+21x_3=17 5(89/9)+8(11/9)+21(-5/6) ~ 41.72... ≠ 17 (not right) so i have made mistakes in the working somewhere

OpenStudy (unklerhaukus):

i dont want to do this any more

OpenStudy (anonymous):

I figured it out a different way. Thank you for your help though.

OpenStudy (unklerhaukus):

can you tell me what you got as an answer ?